OpenStudy (anonymous):

Which line given by its equation below contains the points (1, -1) and (3, 5)? -2y -6x = 0 2y = 6x - 8 y = 3x + 4 y = -3x + 4

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

find the slope first by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (1,-1)...x1 = 1 and y1 = -1 (3,5)...x2 = 3 and y2 = 5 now we sub and find the slope slope(m) = (5 - (-1) / (3 - 1) slope(m) = (5 + 1) / (3 - 1) now finish and tell me the slope

OpenStudy (anonymous):

B. 2y = 6x - 8 @kelliegirl33

OpenStudy (anonymous):

slope(m) = 3 now we use y = mx + b form slope(m) = 3 use either of your points...(1,-1)...x = 1 and y = -1 now we sub -1 = 3(1) + b -1 = 3 + b -1 - 3 = b -4 = b so your equation is : y = 3x - 4.....not there....multiply equation by 2 2y = 6x - 8......thats it....your right

OpenStudy (anonymous):

yea thx :)

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

:D